OpenStudy (anonymous):

What is the polar coordinates of the cartersian points (7,–9)? Other: x^2+y^2=r^2 tan(theta)=y/x

OpenStudy (anonymous):

you got it. make the replacements to see what you get for the number \[r=\sqrt{x^2+y^2}\] and \[\theta=\tan^{-1}(-\frac{9}{7})\]

OpenStudy (anonymous):

Satellite, I have tried the same method to the cartesian coord. of -8,-5, providing me with a r value of square root of 89 (~9.43398), however, the answer does not match up. I am wondering if there is another method.

OpenStudy (anonymous):

oh i see maybe the problem (-8,-5) is in quadrant 3, so you cannot use \[\theta=\tan^{-1}(\frac{y}{x})\] you have to adjust

OpenStudy (anonymous):

you have the right r for sure. i am thinking maybe your theta was off.